TANGENCIES FOR REAL AND COMPLEX HÉNON

MAPS: AN ANALYTIC METHOD

JOHN ERIK FORNÆSS AND ESTELA A. GAVOSTO

Abstract. We present a method for computing generic homoclinic tangencies in the complex Hénon map, based on analytic parametrizations of the stable and unstable manifold, and we discuss applications and consequences of the existence of such tangencies.